KellyAnne Public Relations just paid an annual dividend of $1.27 on its common stock and increases its dividend by 3.4 percent annually. What is the rate of return on this stock if the current stock price is $38.56 a share? Multiple Choice 6.81 percent 7.87 percent 7.04 percent 7.69 percent 7.82 percent
Answer #1

Ans 6.81 percent

P0 = Price of Share
D1 = Current Dividend
Ke = Cost of Equity
g = growth rate
P0 = D1 / (Ke - g)
38.56 = 1.3132 / (Ke - 3.4%)
Ke - 3.4% = 1.3132/ 38.56
Ke - 3.4% = 3.41%
Ke = 6.81%
D1 = D0* (1 + g)
D1 = 1.27* (1 + 3.4%)
D1 = 1.3132